# Asymptotic stability equals exponential stability—while you twist your eyes

### v2

@@ 2   Further comments on related work
Recently, we showed that when $0\in\mathbb{R}^{n}$ is GAS under a continuous vector field $X$, and if this can be asserted using a $C^{1}$ convex Lyapunov function, then, $X$ is straight-line homotopic to $-\partial_{x}$, such that the origin remains GAS along the homotopy [ref:JongeneelSchwan2024TAC]. This is a convenient result, but not a general one.
@@
Unfortunately, for (2.1), $0\in\mathbb{R}^{n}$ is not known to be GAS along the path $[0,1]\ni s\mapsto H(s,x)$. The scalar and linear cases can be understood, however.
@@
As for any $t\in\mathbb{R}_{>0}$, the only fixed point of $\varphi^{t}(\cdot;X)$ is 00, it follows that for $n=1$, the homotopy (2.1) preserves stability since the sign of $x\mapsto H(s,x)$ cannot flip, for otherwise, $\varphi^{s/(1-s)}(x;X)=x$ must hold for some $s\in(0,1)$ and $x\neq 0$. $\circ$
@@ 3   Stability preserving homotopies
Our vector fields are locally essentially bounded, possibly set-valued at 00 and locally Lipschitz on $\mathbb{R}^{n}\setminus\{0\}$.
@@
    for any $t\geq 0$ we have that the time-$t$ map of the semiflow generated by $\mathcal{F}[X]$, that is, $\varphi^{t}(\cdot;\mathcal{F}[X])$, is homotopic to $\varphi^{t}(\cdot;-\partial_{x})$, along time-$t$ maps corresponding to semiflows such that 00 is strongly GAS; and
@@
    in particular, the semiflow $\varphi(\cdot;\mathcal{F}[X])$ is homotopic to $\varphi(\cdot;-\partial_{x})$, along semiflows that preserve 00 being strongly GAS.
@@
The proof proceeds in 5 steps. First we show that the time-$t$ map $\varphi^{t}(\cdot;\mathcal{F}[X])$ can be homotoped along semiflows to a gradient flow $\varphi^{t}(\cdot;-\nabla V)$, such that along the homotopy 00 remains strongly GAS. By exploiting symmetry we show in Step (ii) that an extension of [ref:grune1999asymptotic] allows for showing that for any $\gamma\in\mathcal{K}_{\infty}$ there is a $T\in\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$ (not just $T\in\mathrm{Homeo}(\mathbb{R}^{n};\mathbb{R}^{n})$) such that $V(T^{-1}(x))=\gamma(\|x\|_{2})$. Then, it follows that $T$ is homotopic to $\mathrm{id}_{\mathbb{R}^{n}}$ along a path in $\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$. In Step (iii), we use this map $T$ to homotope $\varphi^{t}(\cdot;-\nabla V)$ to a semiflow, denoted $\widetilde{\varphi}$, that has $V\circ T^{-1}$ as its Lyapunov function. Again, 00 remains strongly GAS along the path. Next, we show in Step (iv) that $\widetilde{\varphi}^{t}$ can be homotoped to $\varphi^{t}(\cdot;-\partial_{x})$. We conclude in Step (v).
(i) Since 00 is strongly GAS under $\dot{x}\in\mathcal{F}[X](x)$ and $X$ satisfies Assumption 3.1, there is a $C^{\infty}$ Lyapunov pair $(V,W)$, that certifies stability of $0\in\mathbb{R}^{n}$ [ref:clarke1998asymptotic, Thm. 1.3], under any Filippov solution. Then, construct the homotopy $H:[0,1]\times\mathbb{R}^{n}\setminus\{0\}\to\mathbb{R}^{n}\setminus\{0\}$ through non-vanishing vector fields, defined by $(s,x)\mapsto H(s,x):=(1-s)\mathcal{F}[X](x)-s\nabla V(x)$. As for any $s\in[0,1]$, $(1-s)\mathcal{F}[X]-s\nabla V$ satisfies Assumption 3.1, plus $\langle\nabla V(x),H(s,x)\rangle\leq-(1-s)W(x)-s\|\nabla V(x)\|_{2}^{2}$ $\forall\,x\in\mathbb{R}^{n}\setminus\{0\}$, we find that 00 is strongly GAS, in the sense of Filippov, under $\dot{x}\in(1-s)\mathcal{F}[X](x)-s\nabla V(x)$, for any fixed $s\in[0,1]$. To construct a homotopy on the level of semiflows, we follow Hartman [hartman2002ordinary, p. 93] and consider the following extended system:
@@
As $V$ has compact sublevel sets (see Definition 1.3), any Filippov solution to $\Sigma_{H}$ is defined for all $t\geq 0$, e.g., see [ref:khalil2002nonlinear, Thm. 3.3]. Now since $(s,x)\mapsto H(s,x)$ is locally Lipschitz away from 00, we can appeal to the Picard-Lindelöf theorem [ref:teschl2012ordinary, Thm. 2.2] and hence, by strong asymptotic stability [ref:clarke1998asymptotic, Def. 2.1] cf. Example 1.4, any solution to $\Sigma_{H}$ is also unique, which allows us to define the time-$t$ map $(s,x)\mapsto\varphi^{t}((s,x);\Sigma_{H})$ for any $t\geq 0$. In its turn, by considering the path $[0,1]\ni s\mapsto\varphi^{t}((s,\cdot);\Sigma_{H})$, we see that this time-$t$ map $\varphi^{t}$ defines a homotopy from the semiflow under $\mathcal{F}[X]$ to the (semi)flow under $-\nabla V$. Specifically, we have that $\varphi^{t}((0,x);\Sigma_{H})=(0,\varphi^{t}(x;\mathcal{F}[X]))$, $\varphi^{t}((1,x);\Sigma_{H})=(1,\varphi^{t}(x;-\nabla V))$ and $\varphi^{t}((s,0);\Sigma_{H})=(s,0)$ $\forall s\in[0,1]$. Now, for any fixed $s\in[0,1]$, overload notation and define the time-$t$ map $\varphi^{t}(\cdot;(s,\Sigma_{H})):=\pi_{2:n+1}\circ\varphi^{t}((s,\cdot);\Sigma_{H})$, for $\pi_{2:n+1}$ the projection on the last $n$ coordinates. It follows that this map checks out as a semiflow, since $\pi_{2:n+1}$ is continuous and for any $x\in\mathbb{R}^{n}$ we have that
@@
(iii) First, we can always pick the homotopy $\widetilde{H}:[0,1]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ from $\mathrm{id}_{\mathbb{R}^{n}}$ to $T^{-1}$ such that 00 is mapped to 00 along the path in $\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$, e.g., for a homeomorphism $K:\mathbb{R}^{n}\to\mathbb{R}^{n}$ such that $K(0)\neq 0$, consider the map $x\mapsto L(x):=K(x)-K(0)$, with $y\mapsto L^{-1}(y)=K^{-1}(y+K(0))$. Now recall that $0\in\mathbb{R}^{n}$ is GAS under $\varphi^{t}(\cdot;-\nabla V)$, i.e., $V(\varphi^{t}(x;-\nabla V))-V(x)\leq-\int^{t}_{0}W_{V}(\varphi^{\tau}(x;-\nabla V))\mathrm{d}\tau$ $\forall(t,x)\in\mathbb{R}_{>0}\times\mathbb{R}^{n}\setminus\{0\}$, with $x\mapsto W_{V}(x):=\|\nabla V(x)\|_{2}^{2}$. Since $\widetilde{H}(s,\cdot)$ is a homeomorphism with $\widetilde{H}(s,0)=0$, we also have
@@
Note that $\widetilde{\varphi}^{t}(\cdot;0)=\varphi^{t}(\cdot;-\nabla V)$ whereas $\widetilde{\varphi}^{t}(\cdot;1)=T\circ\varphi^{t}(\cdot;-\nabla V)\circ T^{-1}$. Also note that continuity of $[0,1]\ni s\mapsto\widetilde{\varphi}^{t}(\cdot;s)$, in particular, continuity of $[0,1]\ni s\mapsto\widetilde{H}(s,\cdot)^{-1}$, follows from $\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$ being a topological group [ref:arens1946topologies] (endowed with the compact-open topology), i.e., combine that the map $\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})\ni h\mapsto h^{-1}$ is continuous with path-connectedness being preserved under a continuous map. Then, by (3.1), we get
@@
Now, using (3.2), define $\widetilde{V}(\cdot;s)\in C^{0}(\mathbb{R}^{n};\mathbb{R}_{\geq 0})$ through $x\mapsto\widetilde{V}(x;s):=V(\widetilde{H}(s,x))$ and similarly, define $x\mapsto\widetilde{W}_{V}(x;s):=W_{V}(\widetilde{H}(s,x))$ Hence, for any $s\in[0,1]$, we have $\widetilde{V}(\widetilde{\varphi}^{t}(x;s);s)-\widetilde{V}(\widetilde{\varphi}^{0}(x;s);s)\leq-\int^{t}_{0}\widetilde{W}_{V}(\widetilde{\varphi}^{\tau}(x;s);s)\mathrm{d}\tau<0$, $\forall(t,x)\in\mathbb{R}_{>0}\times\mathbb{R}^{n}\setminus\{0\}$. Note that $x\mapsto\widetilde{V}(x;0)=V(x)$, $x\mapsto\widetilde{V}(x;1)=V(T^{-1}(x))=\gamma(\|x\|_{2})=:V_{\gamma}(x)$ and $x\mapsto\widetilde{W}_{V}(x;1)=W_{V}(T^{-1}(x))$. Also, as $\widetilde{H}(s,\cdot)$ fixes 00, $\widetilde{V}(\cdot;s)$ will always have the required compact sublevel sets.
@@
In general, $\widetilde{F}$ need not be continuous at 00, this, because $\mathrm{D}T$ need not be continuous at 00. However, $\gamma$ can always be chosen such that $T$ is $C^{1}$ on $\mathbb{R}^{n}$ with $\mathrm{D}T(0)=0$ [ref:grune1999asymptotic, Prop. 1]. Unfortunately, $V_{\gamma}$ is also by no means smooth at 00 for any choice of $\gamma$. In fact, the appropriate $\gamma\in\mathcal{K}_{\infty}$ to guarantee that $T$ is sufficiently regular is of the following form. Let $\alpha\in\mathcal{K}$ be smooth, define $h(r):=\int^{r}_{0}a(\tau)\mathrm{d}\tau$ and set $\gamma:=h^{-1}$. It follows that $\gamma^{\prime}(r)=1/\alpha(\gamma(r))$ is smooth on $(0,+\infty)$, yet, $\lim_{r\to 0^{+}}\gamma^{\prime}(r)=+\infty$, e.g., $r\mapsto\gamma(r)=r^{1/2}$. However, see that on $\mathbb{R}^{n}\setminus\{0\}$ we have $\langle\nabla V_{\gamma}(x),\widetilde{F}(x)\rangle\leq-\widetilde{W}_{V}(x)$, that is,
@@
Then, as in Step (i), construct a homotopy through non-vanishing vector fields on $\mathbb{R}^{n}\setminus\{0\}$, in this case from $\widetilde{F}$ to $-\nabla V_{q}=-\partial_{x}$, with $V_{q}$ being a Lyapunov function asserting stability of 00 throughout. Then, we show there is homotopy from $\varphi^{t}(\cdot;\widetilde{F})$ to $\varphi^{t}(\cdot;-\partial_{x})$ along semiflows, such that 00 is GAS.
(v) Since we have constructed the desired (i.e., preserving stability) homotopy between any time-$t$ map $\varphi^{t}(\cdot;\mathcal{F}[X])$ to $\varphi^{t}(\cdot;-\partial_{x})$, we just observe that all these maps are continuous in $t$ so that we can conclude on the existence of the homotopy from the semiflow $\varphi(\cdot;\mathcal{F}[X])$ to $\varphi(\cdot;-\partial_{x})$. ∎
@@
It is known that the polynomial ODE $\dot{x}=X_{1}(x):=(-x_{1}+x_{1}x_{2},-x_{2})$ does not admit a polynomial Lyapunov function asserting 00 is GAS [ref:ahmadi2011globally]. A non-polynomial Lyapunov function, however, is given by $(x_{1},x_{2})\mapsto V_{1}(x_{1},x_{2}):=\tfrac{1}{2}\log(1+x_{1}^{2})+\tfrac{1}{2}x_{2}^{2}$. To construct a homotopy from $\varphi(\cdot;X_{1})$ to $\varphi(\cdot;-\partial_{x})$, first construct the straight-line homotopy from $X_{1}$ to $-\nabla V_{1}$, with stability asserted throughout by $V_{1}$. Next, define $V(\cdot;s):=(1-s)V_{1}+sV_{q}$, for $s\in[0,1]$ and $x\mapsto V_{q}(x):=\tfrac{1}{2}\|x\|_{2}^{2}$. As for any $s\in[0,1]$, $\nabla V(x;s)=0\iff x=0$, $[0,1]\ni s\mapsto-\nabla V(\cdot;s)$ parametrizes a homotopy from $-\nabla V_{1}$ to $-\partial_{x}$, along vector fields that render 00 GAS, asserted by $V(\cdot;s)$. Although $X_{1}$ does not admit a convex polynomial Lyapunov function, the negative gradient flow of the Lyapunov function corresponding to $X_{1}$ does cf[ref:JongeneelSchwan2024TAC]. $\circ$
@@
$[0,1]\times\mathbb{R}\ni(s,x)\mapsto H_{i}(s,x):=\mathrm{sgn}(x)\left(\frac{1+s}{2}x^{2}+(1-s)\frac{3}{2}\sin(x)^{2}\right)^{2-(3/2)s},$
with $H_{i}(0,\cdot)=T$ and $H_{i}(1,\cdot)=\mathrm{id}_{\mathbb{R}}$ (e.g., $v_{i}=\sqrt{|H_{i}(0,\cdot)|}$). Next, consider the homotopy $(s^{\prime},x)\mapsto x^{1/2+(3/2)s^{\prime}}$ to construct the path from $v_{i}$ to $x\mapsto v_{q}(x):=x^{2}$, along continuous functions such that 00 is the global minimizer throughout [^7] For a simulation of this homotopy, see wjongeneel.nl/figinvex.gif., see Figure 3.1. For functions as simple as $v_{i}$ one can find simpler homotopies (e.g., a straight-line), however, to the best of our knowledge, being able to guarantee the mere existence of such a homotopy is new. Proposition 3.2 provides us with the existence of a homotopy from a smooth Lyapunov function $V$ to $V_{q}$, along continuous [^8] It is not evident, and currently unknown, whether smoothness can be preserved throughout the homotopy, see Step (iii) of the proof of Proposition 3.2. Lyapunov functions that assert $0\in\mathbb{R}^{n}$ is GAS along the homotopy. Differently put, we can find a homotopy from a coercive, invex function, to a convex function, such that along the homotopy the minimizer is preserved. This might be of independent interest. $\circ$
@@
First, since $p^{\star}$ is GAS under $X$, there is $V\in C^{\infty}(M^{n};\mathbb{R}_{\geq 0})$ such that $X(V)=L_{X}V<0$ on $M^{n}\setminus\{p^{\star}\}$ [ref:fathi2019smoothing]. Now fix some Riemannian metric, which exists [Lee2, Ch. 13], and define the Riemannian gradient $\mathrm{grad}\,V$. It follows that $-\mathrm{grad}\,V(V)<0$ on $M^{n}\setminus\{p^{\star}\}$. Then, for any $t\in\mathbb{R}$, the corresponding time-$t$ map $p\mapsto\varphi^{t}(p;-\mathrm{grad}\,V)$ is a diffeomorphism. This implies there is a diffeomorphism $\psi:M^{n}\to\mathbb{R}^{n}$ [ref:wilson1967structure, Thm. 2.2]. Next, define the diffeomorphism $\widetilde{\psi}$ by ${M}^{n}\ni p\mapsto\widetilde{\psi}(p):=\psi(p)-\psi(p^{\star})\in\mathbb{R}^{n}$, such that $\widetilde{\psi}(p^{\star})=0$. Now set $x:=\widetilde{\psi}(p)$ and consider

### v3

@@ 2  Further comments on related work
Recently, we showed that when $0\in\mathbb{R}^{n}$ is GAS under a continuous vector field $X$, and if this can be asserted using a $C^{1}$ convex Lyapunov function, then, $X$ is straight-line homotopic to $\partial_{x}$, such that the origin remains GAS along the homotopy [ref:JongeneelSchwan2024TAC]. This is a convenient result, but not a general one.
@@
Unfortunately, for (2.1), $0\in\mathbb{R}^{n}$ is not known to be GAS along the path $0,1]\ni s\mapsto H(s,x)$. The scalar and linear cases can be understood, however.
@@
As for any $t\in\mathbb{R}_{>0}$, the only fixed point of $\varphi^{t}(\cdot;X)$ is 0, it follows that for $n=1$, the homotopy (2.1) preserves stability since the sign of $x\mapsto H(s,x)$ cannot flip, for otherwise, $\varphi^{s/(1-s)}(x;X)=x$ must hold for some $s\in(0,1)$ and $x\neq 0$.  $\circ$
@@ 3  Stability preserving homotopies
Our vector fields are locally essentially bounded, possibly set-valued at 0 and locally Lipschitz on $\mathbb{R}^{n}\setminus\{0\}$.
@@
    for any $t\geq 0$ we have that the time-$t$ map of the semiflow generated by $\mathcal{F}[X]$, that is, $\varphi^{t}(\cdot;\mathcal{F}[X])$, is homotopic to $\varphi^{t}(\cdot;-\partial_{x})$, along time-$t$ maps corresponding to semiflows such that 0 is strongly GAS; and
@@
    in particular, the semiflow $\varphi(\cdot;\mathcal{F}[X])$ is homotopic to $\varphi(\cdot;-\partial_{x})$, along semiflows that preserve 0 being strongly GAS.
@@
The proof proceeds in 5 steps. First we show that the time-$t$ map $\varphi^{t}(\cdot;\mathcal{F}[X])$ can be homotoped along semiflows to a gradient flow $\varphi^{t}(\cdot;-\nabla V)$, such that along the homotopy 0 remains strongly GAS. By exploiting symmetry we show in Step (ii) that an extension of [ref:grune1999asymptotic] allows for showing that for any $\gamma\in\mathcal{K}_{\infty}$ there is a $T\in\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$ (not just $T\in\mathrm{Homeo}(\mathbb{R}^{n};\mathbb{R}^{n})$) such that $V(T^{-1}(x))=\gamma(\|x\|_{2})$. Then, it follows that $T$ is homotopic to $\mathrm{id}_{\mathbb{R}^{n}}$ along a path in $\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$. In Step (iii), we use this map $T$ to homotope $\varphi^{t}(\cdot;-\nabla V)$ to a semiflow, denoted $\widetilde{\varphi}$, that has $V\circ T^{-1}$ as its Lyapunov function. Again, 0 remains strongly GAS along the path. Next, we show in Step (iv) that $\widetilde{\varphi}^{t}$ can be homotoped to $\varphi^{t}(\cdot;-\partial_{x})$. We conclude in Step (v).
(i) Since 0 is strongly GAS under $\dot{x}\in\mathcal{F}[X](x)$ and $X$ satisfies Assumption 3.1, there is a $C^{\infty}$ Lyapunov pair $(V,W)$, that certifies stability of $0\in\mathbb{R}^{n}$ [ref:clarke1998asymptotic, Thm. 1.3], under any Filippov solution. Then, construct the homotopy $H:[0,1]\times\mathbb{R}^{n}\setminus\{0\}\to\mathbb{R}^{n}\setminus\{0\}$ through non-vanishing vector fields, defined by $(s,x)\mapsto H(s,x):=(1-s)\mathcal{F}[X](x)-s\nabla V(x)$. As for any $s\in[0,1]$, $(1-s)\mathcal{F}[X]-s\nabla V$ satisfies Assumption 3.1, plus $\langle\nabla V(x),H(s,x)\rangle\leq-(1-s)W(x)-s\|\nabla V(x)\|_{2}^{2}$ $\forall\,x\in\mathbb{R}^{n}\setminus\{0\}$, we find that 0 is strongly GAS, in the sense of Filippov, under $\dot{x}\in(1-s)\mathcal{F}[X](x)-s\nabla V(x)$, for any fixed $s\in[0,1]$. To construct a homotopy on the level of semiflows, we follow Hartman [hartman2002ordinary, p. 93] and consider the following extended system:
@@
As $V$ has compact sublevel sets (see Definition 1.3), any Filippov solution to $\Sigma_{H}$ is defined for all $t\geq 0$, e.g., see [ref:khalil2002nonlinear, Thm. 3.3]. Now since $(s,x)\mapsto H(s,x)$ is locally Lipschitz away from 0, we can appeal to the Picard-Lindelöf theorem [ref:teschl2012ordinary, Thm. 2.2] and hence, by strong asymptotic stability [ref:clarke1998asymptotic, Def. 2.1] cf. Example 1.4, any solution to $\Sigma_{H}$ is also unique, which allows us to define the time-$t$ map $(s,x)\mapsto\varphi^{t}((s,x);\Sigma_{H})$ for any $t\geq 0$. In its turn, by considering the path $0,1]\ni s\mapsto\varphi^{t}((s,\cdot);\Sigma_{H})$, we see that this time-$t$ map $\varphi^{t}$ defines a homotopy from the semiflow under $\mathcal{F}[X]$ to the (semi)flow under $\nabla V$. Specifically, we have that $\varphi^{t}((0,x);\Sigma_{H})=(0,\varphi^{t}(x;\mathcal{F}[X]))$, $\varphi^{t}((1,x);\Sigma_{H})=(1,\varphi^{t}(x;-\nabla V))$ and $\varphi^{t}((s,0);\Sigma_{H})=(s,0)$ $\forall s\in[0,1]$. Now, for any fixed $s\in[0,1]$, overload notation and define the time-$t$ map $\varphi^{t}(\cdot;(s,\Sigma_{H})):=\pi_{2:n+1}\circ\varphi^{t}((s,\cdot);\Sigma_{H})$, for $\pi_{2:n+1}$ the projection on the last $n$ coordinates. It follows that this map checks out as a semiflow, since $\pi_{2:n+1}$ is continuous and for any $x\in\mathbb{R}^{n}$ we have that
@@
(iii) First, we can always pick the homotopy $\widetilde{H}:[0,1]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ from $\mathrm{id}_{\mathbb{R}^{n}}$ to $T^{-1}$ such that 0 is mapped to 0 along the path in $\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$, e.g., for a homeomorphism $K:\mathbb{R}^{n}\to\mathbb{R}^{n}$ such that $K(0)\neq 0$, consider the map $x\mapsto L(x):=K(x)-K(0)$, with $y\mapsto L^{-1}(y)=K^{-1}(y+K(0))$. Now recall that $0\in\mathbb{R}^{n}$ is GAS under $\varphi^{t}(\cdot;-\nabla V)$, i.e., $V(\varphi^{t}(x;-\nabla V))-V(x)\leq-\int^{t}_{0}W_{V}(\varphi^{\tau}(x;-\nabla V))\mathrm{d}\tau$ $\forall(t,x)\in\mathbb{R}_{>0}\times\mathbb{R}^{n}\setminus\{0\}$, with $x\mapsto W_{V}(x):=\|\nabla V(x)\|_{2}^{2}$. Since $\widetilde{H}(s,\cdot)$ is a homeomorphism with $\widetilde{H}(s,0)=0$, we also have
@@
Note that $\widetilde{\varphi}^{t}(\cdot;0)=\varphi^{t}(\cdot;-\nabla V)$ whereas $\widetilde{\varphi}^{t}(\cdot;1)=T\circ\varphi^{t}(\cdot;-\nabla V)\circ T^{-1}$. Also note that continuity of $0,1]\ni s\mapsto\widetilde{\varphi}^{t}(\cdot;s)$, in particular, continuity of $0,1]\ni s\mapsto\widetilde{H}(s,\cdot)^{-1}$, follows from $\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})$ being a topological group [ref:arens1946topologies] (endowed with the compact-open topology), i.e., combine that the map $\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})\ni h\mapsto h^{-1}$ is continuous with path-connectedness being preserved under a continuous map. Then, by (3.1), we get
@@
Now, using (3.2), define $\widetilde{V}(\cdot;s)\in C^{0}(\mathbb{R}^{n};\mathbb{R}_{\geq 0})$ through $x\mapsto\widetilde{V}(x;s):=V(\widetilde{H}(s,x))$ and similarly, define $x\mapsto\widetilde{W}_{V}(x;s):=W_{V}(\widetilde{H}(s,x))$ Hence, for any $s\in[0,1]$, we have $\widetilde{V}(\widetilde{\varphi}^{t}(x;s);s)-\widetilde{V}(\widetilde{\varphi}^{0}(x;s);s)\leq-\int^{t}_{0}\widetilde{W}_{V}(\widetilde{\varphi}^{\tau}(x;s);s)\mathrm{d}\tau<0$, $\forall(t,x)\in\mathbb{R}_{>0}\times\mathbb{R}^{n}\setminus\{0\}$. Note that $x\mapsto\widetilde{V}(x;0)=V(x)$, $x\mapsto\widetilde{V}(x;1)=V(T^{-1}(x))=\gamma(\|x\|_{2})=:V_{\gamma}(x)$ and $x\mapsto\widetilde{W}_{V}(x;1)=W_{V}(T^{-1}(x))$. Also, as $\widetilde{H}(s,\cdot)$ fixes 0, $\widetilde{V}(\cdot;s)$ will always have the required compact sublevel sets.
@@
In general, $\widetilde{F}$ need not be continuous at 0, this, because $\mathrm{D}T$ need not be continuous at 0. However, $\gamma$ can always be chosen such that $T$ is $C^{1}$ on $\mathbb{R}^{n}$ with $\mathrm{D}T(0)=0$ [ref:grune1999asymptotic, Prop. 1]. Unfortunately, $V_{\gamma}$ is also by no means smooth at 0 for any choice of $\gamma$. In fact, the appropriate $\gamma\in\mathcal{K}_{\infty}$ to guarantee that $T$ is sufficiently regular is of the following form. Let $\alpha\in\mathcal{K}$ be smooth, define $h(r):=\int^{r}_{0}a(\tau)\mathrm{d}\tau$ and set $\gamma:=h^{-1}$. It follows that $\gamma^{\prime}(r)=1/\alpha(\gamma(r))$ is smooth on $(0,+\infty)$, yet, $\lim_{r\to 0^{+}}\gamma^{\prime}(r)=+\infty$, e.g., $r\mapsto\gamma(r)=r^{1/2}$. However, see that on $\mathbb{R}^{n}\setminus\{0\}$ we have $\langle\nabla V_{\gamma}(x),\widetilde{F}(x)\rangle\leq-\widetilde{W}_{V}(x)$, that is,
@@
[…]
(v) Since we have constructed the desired (i.e., preserving stability) homotopy between any time-$t$ map $\varphi^{t}(\cdot;\mathcal{F}[X])$ to $\varphi^{t}(\cdot;-\partial_{x})$, we observe that all these maps are continuous in $t$, in fact, all steps show that joint continuity holds so that we can conclude on the existence of the homotopy from the semiflow $\varphi(\cdot;\mathcal{F}[X])$ to $\varphi(\cdot;-\partial_{x})$. ∎
@@
It is known that the polynomial ODE $\dot{x}=X_{1}(x):=(-x_{1}+x_{1}x_{2},-x_{2})$ does not admit a polynomial Lyapunov function asserting 0 is GAS [ref:ahmadi2011globally]. A non-polynomial Lyapunov function, however, is given by $(x_{1},x_{2})\mapsto V_{1}(x_{1},x_{2}):=\tfrac{1}{2}\log(1+x_{1}^{2})+\tfrac{1}{2}x_{2}^{2}$. To construct a homotopy from $\varphi(\cdot;X_{1})$ to $\varphi(\cdot;-\partial_{x})$, first construct the straight-line homotopy from $X_{1}$ to $\nabla V_{1}$, with stability asserted throughout by $V_{1}$. Next, define $V(\cdot;s):=(1-s)V_{1}+sV_{q}$, for $s\in[0,1]$ and $x\mapsto V_{q}(x):=\tfrac{1}{2}\|x\|_{2}^{2}$. As for any $s\in[0,1]$, $\nabla V(x;s)=0\iff x=0$, $0,1]\ni s\mapsto-\nabla V(\cdot;s)$ parametrizes a homotopy from $\nabla V_{1}$ to $\partial_{x}$, along vector fields that render 0 GAS, asserted by $V(\cdot;s)$. Although $X_{1}$ does not admit a convex polynomial Lyapunov function, the negative gradient flow of the Lyapunov function corresponding to $X_{1}$ does cf. [ref:JongeneelSchwan2024TAC].  $\circ$
@@
$0,1]\times\mathbb{R}\ni(s,x)\mapsto H_{i}(s,x):=\mathrm{sgn}(x)\left(\frac{1+s}{2}x^{2}+(1-s)\frac{3}{2}\sin(x)^{2}\right)^{2-(3/2)s},$
with $H_{i}(0,\cdot)=T$ and $H_{i}(1,\cdot)=\mathrm{id}_{\mathbb{R}}$ (e.g., $v_{i}=\sqrt{|H_{i}(0,\cdot)|}$). Next, consider the homotopy $(s^{\prime},x)\mapsto x^{1/2+(3/2)s^{\prime}}$ to construct the path from $v_{i}$ to $x\mapsto v_{q}(x):=x^{2}$, along continuous functions such that 0 is the global minimizer throughout [^7] For a simulation of this homotopy, see wjongeneel.nl/figinvex.gif., see Figure 3.1. For functions as simple as $v_{i}$ one can find simpler homotopies (e.g., a straight-line), however, to the best of our knowledge, being able to guarantee the mere existence of such a homotopy is new. Proposition 3.2 provides us with the existence of a homotopy from a smooth Lyapunov function $V$ to $V_{q}$, along continuous [^8] It is not evident, and currently unknown, whether smoothness can be preserved throughout the homotopy, see Step (iii) of the proof of Proposition 3.2. Lyapunov functions that assert $0\in\mathbb{R}^{n}$ is GAS along the homotopy. Differently put, we can find a homotopy from a coercive, invex function, to a convex function, such that along the homotopy the minimizer is preserved. This might be of independent interest.  $\circ$
@@
First, since $p^{\star}$ is GAS under $X$, there is $V\in C^{\infty}(M^{n};\mathbb{R}_{\geq 0})$ such that $X(V)=L_{X}V<0$ on $M^{n}\setminus\{p^{\star}\}$ [ref:fathi2019smoothing]. Now fix some Riemannian metric, which exists [Lee2, Ch. 13], and define the Riemannian gradient $\mathrm{grad}\,V$. It follows that $\mathrm{grad}\,V(V)<0$ on $M^{n}\setminus\{p^{\star}\}$. Then, for any $t\in\mathbb{R}$, the corresponding time-$t$ map $p\mapsto\varphi^{t}(p;-\mathrm{grad}\,V)$ is a diffeomorphism. This implies there is a diffeomorphism $\psi:M^{n}\to\mathbb{R}^{n}$ [ref:wilson1967structure, Thm. 2.2]. Next, define the diffeomorphism $\widetilde{\psi}$ by ${M}^{n}\ni p\mapsto\widetilde{\psi}(p):=\psi(p)-\psi(p^{\star})\in\mathbb{R}^{n}$, such that $\widetilde{\psi}(p^{\star})=0$. Now set $x:=\widetilde{\psi}(p)$ and consider